If the frequency of an electromagnetic wave is
1 × 10^6 Hz what is the wavelength?
Answer in units of nm
1
Expert's answer
2012-09-07T10:35:33-0400
Let's use the following formula:
L = c / f,
where c is the speed of the
light in vacuum and f is the frequency.
We get the following:
L = 3 * 10^8
/ 10^6 = 300 m = 3 * 10^11 nm.